OpenStudy (anonymous):

Can you help mt find the solutions? y= x²-2x-2 y+4x+5

OpenStudy (anonymous):

no i dont im sorry.. :(

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (anonymous):

@amistre64 can you hilp? i factored it down to x²-2x+3 but i dont know what 2 do next

OpenStudy (amistre64):

y= x^2-2x-2 y=4x+5 since y = y, rewrite this as x^2-2x-2 = 4x+5 , and move it all to one side x^2-6x-7 = 0 is hat i beleive you are "stuck" at

OpenStudy (anonymous):

ok so what do you do next?

OpenStudy (amistre64):

well, its good to understand that multiplying is far simpler than factoring. if this is going to factor nicely, it will have to take the form: (x+a)(x+b) can you expand that out for me?

OpenStudy (anonymous):

is that the same as foiling?

OpenStudy (amistre64):

yes, "FOIL" is a memory device used for people who forget how to multiply things together. So if you must "FOIL", then by all means do so

OpenStudy (amistre64):

when we "foil" that general setup out, we will have something to compare our x^2 -6x -7 to

OpenStudy (anonymous):

x²+xb+ax+ab

OpenStudy (amistre64):

good; lets clean that up a little by combining like terms and compare it to what we have already x^2 +(a+b)x + (ab) x^2 + (-6) x + (-7) we see that the sum of a and b should equal -6, and the product of a and b should equal -7 a+b = -6 ab = -7 this is the formal setup for factoring this into (x+a)(x+b)

OpenStudy (amistre64):

the idea most books have it to do a trial and error. determine some integer factors of -7, and see if they add up to -6

OpenStudy (anonymous):

1 -7?

OpenStudy (amistre64):

that would work yes :) so let a = -7, and b=1 (x+a)(x+b) becomes (x-7)(x+1)

OpenStudy (amistre64):

now, we want to know what this factored form is equal to zero; so solve for x-7 = 0 x+1 = 0 and those will be the x values for the system. the y values are simply found using one of the equation that was originally given, the linear one is simplest to solve. y = 4x+5, for each given value of x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so the answer is -1, 1 and 7,33?